# QuantFace: Low-Bit Post-Training Quantization for One-Step Diffusion Face Restoration

### v1

@@ Abstract
Diffusion models have been achieving remarkable performance in face restoration. However, the heavy computations of diffusion models make it difficult to deploy them on devices like smartphones. In this work, we propose QuantFace, a novel low-bit quantization for one-step diffusion face restoration models, where the full-precision (i.e., 32-bit) weights and activations are quantized to 4$\sim$6-bit. We first analyze the data distribution within activations and find that they are highly variant. To preserve the original data information, we employ rotation-scaling channel balancing. Furthermore, we propose Quantization-Distillation Low-Rank Adaptation (QD-LoRA) that jointly optimizes for quantization and distillation performance. Finally, we propose an adaptive bit-width allocation strategy. We formulate such a strategy as an integer programming problem, which combines quantization error and perceptual metrics to find a satisfactory resource allocation. Extensive experiments on the synthetic and real-world datasets demonstrate the effectiveness of QuantFace under 6-bit and 4-bit. QuantFace achieves significant advantages over recent leading low-bit quantization methods for face restoration. The code is available at https://github.com/jiatongli2024/QuantFace.
@@ 1 Introduction
Face restoration seeks to recover high-quality (HQ) facial images from degraded low-quality (LQ) inputs. The LQ inputs are generated from complex degradation processes, such as blurring, noise, downsampling, and JPEG compression. With the rapid development of deep generative models, generative adversarial networks (GANs) [1] and diffusion-based methods [2, 3, 4] have attracted increasing attention. Several approaches [5, 6, 7, 8] have already demonstrated impressive capability in restoring HQ human faces. Nevertheless, GANs typically suffer from training instability due to adversarial learning, while multi-step diffusion models demand extensive computational resources.
@@
To accelerate inference, recent studies have proposed one-step diffusion (OSD) [9, 10, 11] schemes that reduce the denoising step to one. At the same time, these models largely preserve the performance of the original multi-step diffusion models. Thanks to these advanced techniques, one-step diffusion face restoration (OSDFR) models [12] have become feasible. However, despite this progress, deploying such models on resource-constrained platforms (e.g., smartphones or other mobile devices) remains challenging, because they still require prohibitive computational resources. This limitation hinders the widespread adoption of diffusion-based face restoration methods in real-world scenarios.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LQ DiffBIR [13] OSDFace [12] PassionSR [14] SVDQuant [15] QuantFace (ours) # Step / Bits 50 / W32A32 1 / W32A32 1 / W4A4 1 / W4A4 1 / W4A4 Param. (M) / Ops (G) 1,618 /49,056 978 / 4,264 122 / 533 148 / 728 148 / 728
@@
Figure 1: Visual comparison between the multi-step diffusion and one-step diffusion face restoration models in full-precision, recent quantization methods in 4-bit, and our QuantFace in 4-bit. Our method achieves an 84.85% parameter compression and an 82.91% speedup compared with the full-precision OSDFace [12].
@@
Model quantization [16] offers an effective avenue to further accelerate OSDFR models. By mapping both weights and activations from full-precision (FP) to low-bit precision, quantization significantly reduces memory footprint and computational overhead. Nevertheless, the performance gap between the quantized and the FP version is inevitable, especially under the aggressive 4-bit setting. Minimizing this gap is therefore vital for the successful implementation of quantized OSDFR models.
@@
Although current low-bit quantization strategies have achieved promising results for multi-step diffusion models and for text-to-image generation [17, 18, 19, 15, 20, 21], significant performance drops occur when we apply these methods to OSDFR models. There are three main challenges in low-bit quantization for OSDFR. First, the distribution of dynamic activation values is poorly preserved by static quantizers. Second, current distillation schemes from full-precision (FP) to quantized models are suboptimal. Third, the limited computational resources are not properly allocated.
@@
In this work, we propose QuantFace, a unified and effective low-bit quantization framework for OSDFR models. We select OSDFace [12] as our quantization backbone due to its excellent performance. To address the aforementioned challenges, we propose tailored solutions for each. First, we conduct a detailed analysis of the activation distributions in OSDFace and find that different channels of the activation values exhibit high variance. Previous scaling based methods [22, 15] can remove the outliers from activations to weights and thus decrease quantization error. However, these methods are ineffective in preserving the original activation distribution after per-tensor quantization. We find that the distribution information is vital for basic facial structure generation. To solve the problem, we propose rotation-scaling channel balancing. We consider rotation and scaling simultaneously to maintain a uniform activation distribution. Second, to promote alignment between the FP model and the quantized model, we propose Quantization-Distillation Low-Rank Adaptation (QD-LoRA). QD-LoRA effectively compensates for quantization error while facilitating distillation between the FP model and the quantized model. Third, we propose adaptive bit-width allocation. We analyze the layer-wise sensitivity to quantization and configure the bottlenecks. We design an integer programming-based algorithm for mixed-precision activation configuration. This algorithm incorporates perceptual metrics and numeric reconstruction error as the optimization target and promotes reasonable resource allocation.
@@
We have conducted extensive evaluations on both synthetic and real-world datasets. As shown in Fig. 1, under 4-bit quantization, our method achieves a significant speedup compared to the multi-step face restoration model DiffBIR [13] while preserving comparable visual quality. Furthermore, Tab. 2 demonstrates that at 4-bit precision, we attain parameter and computational cost compression ratios of 84.85% and 82.91%, respectively, compared with OSDFace [12]. These findings validate the efficacy of our proposed approaches for quantizing OSDFR models.
@@
  • We conduct a detailed analysis of the activation distribution of the OSDFR model and identify that quantization of activation values is crucial for preserving facial features. We configure that the inadequate processing of highly variant activations is the key reason for the performance degradation in quantization.
@@ 2.1 Face Restoration
Face restoration seeks to reconstruct high-fidelity facial images from degraded low-quality inputs affected by diverse and complex distortions. A central challenge in this domain is how to incorporate facial priors efficiently and effectively. While classical methods have long leveraged statistical priors, recent advances in deep generative techniques(such as GANs [1] and diffusion models [2, 3, 23]) have driven growing interest in generative priors.
In particular, diffusion-based priors [5, 6, 7, 8] have attracted considerable attention. These methods generate HQ images by gradually removing noise from LQ inputs. Concurrently, vector-quantized (VQ) prior approaches [24, 25, 26, 27] are promising. They map LQ features to discrete codebook entries and then decode the HQ images from the codebook.
The combination of VQ priors and generative diffusion priors has become very popular recently. However, while these hybrid methods can generate excellent human face images, the computational demands become severe. Consequently, developing efficient and effective compression techniques for these models has become a major focus of current research.
@@ 2.2 Model Quantization
Model quantization improves computational efficiency by reducing the precision of model parameters while maintaining performance. Quantization methods are broadly classified into two categories based on whether weight retraining is involved: post-training quantization (PTQ) [28, 29] and quantization-aware training (QAT) [30, 31, 32]. Furthermore, quantization has been demonstrated as an effective technique for compressing large language models (LLMs) [33, 34, 35, 22] for deployment on edge devices. With the rapid advancements in diffusion models (DMs), there has been a growing focus on enhancing their efficiency through quantization techniques including PTQ methods such as PTQ4DM [19], Q-Diffusion [36], and PTQD [18], as well as QAT methods like Q-DM [37]. Recently, EfficientDM [21] introduced a low-rank quantization fine-tuning strategy, while SVDQuant [15] employs 16-bit parallel low-rank branches to preserve performance. PassionSR [14] presents an innovative quantization strategy, achieving 8-bit and 6-bit quantization for one-step diffusion-based super-resolution (OSDSR) models. However, research on the quantization of OSDFR models remains limited, despite their significant differences from traditional DMs.
3 Methods
3.1 Preliminaries
Diffusion Model.
Diffusion models [23] employ a two-stage procedure. First, a complex data distribution is gradually corrupted by successive injections of Gaussian noise. Second, a neural network learns the inverse of this noising process to restore the original samples. In the forward diffusion, the original sample $\mathbf{x}_{0}$ is progressively transformed into pure Gaussian noise $\mathbf{x}_{T}$ by injecting random noise over $T$ steps, formally defined as [3]:
where $\bar{\alpha}_{t}$ is a scheduling parameter, and $t$ denotes the time step.
After $T$ iterations, the data pattern converges to a standard Gaussian distribution: $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$. Reversed diffusion process reconstructs the original distribution by predicting the noise $\epsilon_{\theta}(\mathbf{x}_{t},t)$ formulated as follows [3]:
$\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}})+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}(\mathbf{x}_{t},t)+\sigma_{t}\epsilon_{t},$ (2)
where $\epsilon_{\theta}(\mathbf{x}_{t},t)$ represents the output of the noise prediction network, $\sigma^{2}_{t}$ denotes the noise variance at step $t$ ,and $\epsilon_{t}$ is random noise irrelevant to $\mathbf{x}_{t}$. By gradually predicting the random noise at timestep $t$, the model progressively optimizes $\mathbf{x}_{t}$ to $\mathbf{x}_{0}$, and ultimately generates high-quality samples. Our one-step estimation can be formulated as [9]:
Model Quantization.
Model quantization [38] is an effective alternative to reduce memory and computational costs by converting floating-point numbers to fixed-point numbers. This process utilizes approximation with integer $x_{\text{int}}$ and quantization parameters (scaling factor $s$, zero point $z$). The fake quantization process can be defined as [16]:
[…]
To address the non-differentiability of the rounding operation, we employ the straight-through estimator (STE [39]) to approximate the gradients during backpropagation:
3.2 Rotation-Scaling Channel Balancing
Figure 3: Overview of our QuantFace. First, under the 4-bit precision setting, we use the quantization errors and perceptual importance weights as the objective for integer programming, and allocate appropriate precision to activation of each layer. Second, before training, we integrate the scaling factor and rotation matrix into the weights, and only apply an online Randomized Hadamard Transform in the convolution layers. Third, we align the quantized model with the FP model on the calibration dataset by optimizing the dual-branch low-rank matrices we design.
In any quantization method, appropriately managing the relationship between outliers and non-outliers is of critical importance. Existing channel-scaling methods can shift the quantization difficulty from activations to weights [15, 22]. After scaling, activations exhibit reduced magnitude and fewer outliers, resulting in lower quantization error. However, when it comes to convolution neural networks like Unet [40], we typically employ per-tensor quantization for activations, which means a single set of quantization parameters is shared across all channels of a tensor. Although the aforementioned scaling method suppresses extreme outliers, inter-channel imbalance still remains. When the bit-width is extremely low (e.g., 4-bit), the performance of the quantized model drops significantly.
To solve the problem, incoherent processing via Randomized Hadamard Transform (RHT) has recently emerged as a promising technique [41], and it has been shown to be a remarkable quantization method for LLMs [42, 43] and ViTs [44]. This approach requires no parameter tuning and can effectively generate a uniform data distribution across channels, thereby maximally preserving information after quantization. However, this method cannot adequately smooth extreme outliers.
To the best of our knowledge, prior quantization works primarily focused on Transformer architectures. We are the first to design a scaling and rotation channel balancing method for convolution neural networks like UNet [40]. Through our analysis, we show that convolution can be transformed to matrix multiplication to some degree. The original 2D-convolution can be written as:
$y_{f,i,j}=\displaystyle\sum_{u=1}^{K_{w}}\sum_{v=1}^{K_{h}}\sum_{c=1}^{C}k_{f,c,u,v}x_{c,i+u-1,j+v-1}=\displaystyle\sum_{u=1}^{K_{w}}\sum_{v=1}^{K_{h}}\textbf{k}_{f,u,v}^{T}\textbf{x}_{i+u-1,j+v-1},$ (6)
$\textbf{k}_{f,u,v}\triangleq\textbf{k}\in\mathbb{R}^{C},\textbf{x}_{h+u-1,w+v-1}\triangleq\textbf{x}\in\mathbb{R}^{C},$ (7)
so we have:
$\textbf{k}^{T}\textbf{x}=(\textbf{k}^{T}\odot s)(\frac{1}{s}\odot\textbf{x})=\textbf{k}^{T}{\mathbf{H}}^{T}\mathbf{H}\mathbf{x},\quad s=\frac{\underset{(h,w)}{\max}(|\mathbf{X}|)^{\alpha}}{\underset{(C,K_{w},K_{h})}{\max}(|\mathbf{K}|)^{1-\alpha}},$ (8)
where $\mathbf{K}\in\mathbb{R}^{F\times C\times K_{w}\times K_{h}},\mathbf{X}\in\mathbb{R}^{C\times h\times w}$ is convolution kernel and input tensor of the layer. $\mathbf{H}$ is a Hadamard matrix. We find that for convolution layers in OSDFace [12], the corresponding Hadamard matrices exist. According to [41], only a minimal amount of memory is required to store the corresponding random ±$\pm 1$ vector of each matrix.
@@ 3.2 Rotation-Scaling Channel Balancing
Figure 4: A sample data distribution along input channels of unet.down_blocks.1.resnets.1.conv2 activations. They are processed with different balancing techniques. Directly employing the Scaling or Randomized Hadamard Transformation will not result in satisfactory distribution for quantization.

### v2

# QuantFace: Efficient Quantization for Face Restoration

@@ Abstract
[…]
@@ 1 Introduction
Face restoration seeks to recover high-quality (HQ) facial images from degraded low-quality (LQ) inputs. The LQ inputs are generated from complex degradation processes, such as blurring, noise, downsampling, and JPEG compression. With the rapid development of deep generative models, generative adversarial networks (GANs) [goodfellow2014generative] and diffusion-based methods [ddpm, ddim, song2020score] have attracted increasing attention. Several approaches [kawar2022ddrm, wang2022ddnm, fei2022gdp, yang2023pgdiff] have already demonstrated impressive capability in restoring HQ human faces. Nevertheless, GANs typically suffer from training instability due to adversarial learning, while multi-step diffusion models demand extensive computational resources.
@@
To accelerate inference, recent studies have proposed one-step diffusion (OSD) [dmd, song2023consistency, li2024distillation] schemes that reduce the denoising step to one. At the same time, these models largely preserve the performance of the original multi-step diffusion models. Thanks to these advanced techniques, one-step diffusion face restoration (OSDFR) models [osdface] have become feasible. However, despite this progress, deploying such models on resource-constrained platforms (e.g., smartphones or other mobile devices) remains challenging, because they still require prohibitive computational resources. This limitation hinders the widespread adoption of diffusion-based face restoration methods in real-world scenarios.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LQ DiffBIR [lin2024diffbir] OSDFace [osdface] SVDQuant [li2025svdquant] QuantFace (ours) # Step / Bits 50 / 32-bit 1 / 32-bit 1 / 4-bit 1 / 4-bit
@@
Figure 1: Visual comparison between the multi-step diffusion and one-step diffusion face restoration models in full-precision, recent quantization methods in 4-bit, and our QuantFace in 4-bit. Our method achieves an 84.85% parameter compression and an 82.91% speedup compared with the full-precision OSDFace [osdface].
@@
Model quantization [white] offers an effective avenue to further accelerate OSDFR models. By mapping both weights and activations from full-precision (FP) to low-bit precision, quantization significantly reduces memory footprint and computational overhead. Nevertheless, the performance gap between the quantized and the FP version is inevitable, especially under the aggressive 4-bit setting. Minimizing this gap is therefore vital for the successful implementation of quantized OSDFR models.
@@
Although current low-bit quantization strategies have achieved promising results for multi-step diffusion models and for text-to-image generation [qdiffusion, PTQD, PTQ4DM, li2025svdquant, sui2024bitsfusion, He2023EfficientDM], significant performance drops occur when we apply these methods to OSDFR models. There are three main challenges in low-bit quantization for OSDFR. First, the distribution of dynamic activation values is vital for high-frequency facial details but it is poorly preserved by static quantizers. Second, current calibration schemes from FP to quantized models are suboptimal. Third, the limited computational resources are allocated improperly.
@@
In this work, we propose QuantFace, a unified and effective low-bit quantization framework for OSDFR models. We select OSDFace [osdface] as our quantization backbone due to its excellent performance. To address the aforementioned challenges, we propose tailored solutions for each. First, we conduct a detailed analysis of the activation distributions in OSDFace and find that different channels of the activation values exhibit high variance. Previous scaling-based methods [smoothquant, li2025svdquant] can remove the outliers from activations to weights and thus decrease quantization error. However, these methods are ineffective in preserving the original activation distribution after per-tensor quantization. We find that the distribution information is vital for basic facial structure generation. To solve the problem, we propose rotation-scaling channel balancing. We consider rotation and scaling simultaneously to maintain a uniform activation distribution. Second, to promote alignment between the FP model and the quantized model, we propose Quantization-Distillation Low-Rank Adaptation (QD-LoRA). QD-LoRA effectively compensates for quantization error while facilitating distillation between the FP model and the quantized model. Third, we propose adaptive bit-width allocation. We analyze the layer-wise sensitivity to quantization and configure the bottlenecks. We design an integer programming-based algorithm for mixed-precision activation configuration. This algorithm incorporates perceptual metrics and numeric reconstruction error as the optimization target and promotes reasonable resource allocation.
@@
We have conducted extensive evaluations on both synthetic and real-world datasets. As shown in Fig. 1, even under the aggressive 4-bit setting, our model achieves visual quality comparable to the multi-step diffusion model DiffBIR [lin2024diffbir]. Furthermore, Tab. 2 demonstrates that at 4-bit precision, we attain parameter and computational cost compression ratios of 84.85% and 82.91%, respectively, compared with OSDFace [osdface]. These findings validate the efficacy of our proposed approaches for quantizing OSDFR models.
@@
  • We conduct a detailed analysis of the activation distribution of the OSDFace and identify that quantization of activation values is crucial for preserving facial features. We configure that the inadequate processing of highly variant activations is the key reason for the performance degradation in quantization.
@@
Figure 3: Overview of our QuantFace. First, under the 4-bit precision setting, we use the quantization errors and perceptual importance weights as the objective for integer programming, and allocate appropriate precision to activation of each layer. Second, before training, we integrate the scaling factor and rotation matrix into the weights, and only apply an online RHT in the convolution layers. Third, we align the quantized model with the FP model on the calibration dataset by optimizing the dual-branch low-rank matrices we design.
@@ 2.1 Face Restoration
Face restoration seeks to reconstruct high-fidelity facial images from degraded low-quality inputs affected by diverse and complex distortions. A central challenge in this domain is how to incorporate facial priors efficiently and effectively. While classical methods have long leveraged statistical priors, recent advances in deep generative techniques (such as GANs [goodfellow2014generative] and diffusion models [ddpm, ddim, rombach2022highresolutionimagesynthesislatent]) have driven growing interest in generative priors.
In particular, diffusion-based priors [kawar2022ddrm, wang2022ddnm, fei2022gdp, yang2023pgdiff] have attracted considerable attention. These methods generate HQ images by gradually removing noise from LQ inputs. Concurrently, vector-quantized (VQ) prior approaches [vqfr, zhou2022codeformer, wang2023restoreformer++, daefr] are promising. They map LQ features to codebook entries and then decode the HQ images from the codebook.
The combination of VQ priors and generative diffusion priors has become very popular recently. However, while these hybrid methods can generate excellent human face images, the computational demands become severe. Consequently, developing effective compression techniques for these models has become a major focus of current research.
@@ 2.2 Model Quantization
Model quantization improves computational efficiency by reducing the precision of model parameters while maintaining performance. Quantization methods are broadly classified into two categories based on whether weight retraining is involved: post-training quantization (PTQ) [ZeroQuant, brecq] and quantization-aware training (QAT) [Bhalgat2020LSQ+, efficientqat, yu2024improving]. Furthermore, quantization has been demonstrated as an effective technique for compressing large language models (LLMs) [liu2023llm, bondarenko2024low, li2023loftq, smoothquant] for deployment on edge devices. With the rapid advancements in diffusion models (DMs), there has been a growing focus on enhancing their efficiency through quantization techniques including PTQ methods such as PTQ4DM [PTQ4DM], Q-Diffusion [li2023q], and PTQD [PTQD], as well as QAT methods like Q-DM [qdm]. Recently, EfficientDM [He2023EfficientDM] introduced a low-rank quantization fine-tuning strategy, while SVDQuant [li2025svdquant] employs 16-bit parallel low-rank branches to preserve performance. PassionSR [zhu2024passionsr] presents an innovative quantization strategy, achieving 8-bit and 6-bit quantization for one-step diffusion-based super-resolution (OSDSR) models. However, existing works on quantizing image restoration models to 4-bit remain limited.
@@
Figure 4: A sample data distribution along input channels of unet.down_blocks.1.resnets.1.conv2 activations. They are processed with different balancing techniques. Directly employing the Scaling or RHT will not result in satisfactory distribution for quantization.
3 Methods
3.1 Preliminaries
Diffusion Model.
Diffusion models [rombach2022highresolutionimagesynthesislatent] employ a two‐stage procedure. First, a complex data distribution is gradually corrupted by successive injections of Gaussian noise. Second, a neural network learns the inverse of this noising process to restore the original samples. In the forward diffusion, the original sample $\mathbf{x}_{0}$ is progressively transformed into pure Gaussian noise $\mathbf{x}_{T}$ by injecting random noise over $T$ steps, formally defined as [ddim]:
where $\bar{\alpha}_{t}$ is a scheduling parameter, and $t$ is the time step.
After $T$ iterations, the data pattern converges to a standard Gaussian distribution: $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$. Reversed diffusion process reconstructs the original distribution by predicting the noise $\epsilon_{\theta}(\mathbf{x}_{t},t)$ formulated as follows [ddim]:
$\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}})$ (2)
$+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}(\mathbf{x}_{t},t)+\sigma_{t}\epsilon_{t},$
where $\epsilon_{\theta}(\mathbf{x}_{t},t)$ represents the output of the noise prediction network, $\sigma^{2}_{t}$ denotes the noise variance at step $t$, and $\epsilon_{t}$ is random noise irrelevant to $\mathbf{x}_{t}$. By gradually predicting the random noise at timestep $t$, the model progressively optimizes $\mathbf{x}_{t}$ to $\mathbf{x}_{0}$, and ultimately generates high-quality samples. Our one-step estimation [dmd] can be formulated as:
Model Quantization.
Model quantization [jacob2017quantizationtrainingneuralnetworks] is an effective alternative to reduce memory and computational costs by converting floating-point numbers to fixed-point numbers. This process utilizes approximation with integer $x_{\text{int}}$ and quantization parameters (scaling factor $s$, zero point $z$). The fake quantization [white] process can be defined as:
where $x_{\text{int}}$ refers to the quantized integer and $\hat{x}$ is the simulated quantized float-point number. $\left\lfloor\cdot\right\rceil$ is the round-to-nearest operator, and $\text{clamp}(\cdot,l,u)$ ensures values remain within the range $[l,u]$.
To address the non-differentiability of the rounding operation, we employ the straight-through estimator (STE [liu2022nonuniform]) to approximate the gradients during backpropagation:
3.2 Rotation-Scaling Channel Balancing
In any quantization method, appropriately managing the relationship between outliers and non‑outliers is of critical importance. Existing channel-scaling methods can shift the quantization difficulty from activations to weights [li2025svdquant, smoothquant]. After scaling, activations exhibit reduced magnitude and fewer outliers, resulting in lower quantization error. However, when it comes to convolution neural networks like UNet [unet], we typically employ per‑tensor quantization for activations, which means a single set of quantization parameters is shared across all channels of a tensor. Although the aforementioned scaling method suppresses extreme outliers, inter-channel imbalance remains. When the bit‑width is extremely low (e.g., 4-bit), the performance of the quantized model drops significantly.
To solve the problem, incoherent processing via Randomized Hadamard Transform (RHT) has recently emerged as a promising technique [quipsharp], and it is a remarkable quantization method for LLMs [quarot, spinquant] and ViTs [vidit]. This approach requires no parameter tuning and can effectively generate a uniform data distribution across channels, thereby maximally preserving information after quantization. However, this method cannot adequately smooth extreme outliers.
To the best of our knowledge, prior quantization works have primarily focused on Transformer architectures. We are the first to design a scaling and rotation channel balancing method for convolution neural networks like UNet [unet]. Through our analysis, we show that convolution can be transformed to matrix multiplication to some degree. The original 2D-convolution can be written as:
$y_{f,i,j}$ $=\sum_{u=1}^{K_{w}}\sum_{v=1}^{K_{h}}\sum_{c=1}^{C}k_{f,c,u,v}x_{c,i+u-1,j+v-1}$ (6)
$=\sum_{u=1}^{K_{w}}\sum_{v=1}^{K_{h}}\textbf{k}_{f,u,v}^{T}\textbf{x}_{i+u-1,j+v-1},$
$\textbf{k}_{f,u,v}\triangleq\textbf{k}\in\mathbb{R}^{C\times 1},\textbf{x}_{h+u-1,w+v-1}\triangleq\textbf{x}\in\mathbb{R}^{C\times 1},$ (7)
so we have:
$\textbf{k}^{T}\textbf{x}$ $=(\textbf{k}^{T}\odot s)(\frac{1}{s}\odot\textbf{x})=\textbf{k}^{T}{\mathbf{H}}^{T}\mathbf{H}\mathbf{x},$ (8)
$s$ $=\frac{\underset{(h,w)}{\max}(|\mathbf{X}|)^{\alpha}}{\underset{(C,K_{w},K_{h})}{\max}(|\mathbf{K}|)^{1-\alpha}},$
where $\mathbf{K}\in\mathbb{R}^{F\times C\times K_{w}\times K_{h}},\mathbf{X}\in\mathbb{R}^{C\times h\times w}$ is convolution kernel and input tensor of the layer. $\mathbf{H}$ is a Hadamard matrix. We find that for convolution layers in OSDFace [osdface], the corresponding Hadamard matrices exist. According to [quipsharp], only a minimal amount of memory is required to store the corresponding random $\pm 1$ vector of each matrix.